OpenStudy (anonymous):

6. Let X and Y be two r.v.s. with E(X) = E(Y ) = 0, Cov(X, Y ) = 1, V (X) = 1 and V (Y ) = 2. What is the expectation of (X + 1)(X − 3Y )? Need help I know the method but answer is wrong help...

OpenStudy (anonymous):

Well, as a educated guess I know what you're talking about

OpenStudy (anonymous):

this is kind of statistics or probability or something in between.

OpenStudy (anonymous):

\[\begin{align*}E\left[(X+1)(X-3Y)\right]&=E\left[X^2+X-3XY-3Y\right]\\ &=E\left[X^2\right]+E\left[X\right]-3E\left[XY\right]-3E\left[Y\right]\\ &=E\left[X^2\right]+0-3E\left[XY\right]-3(0) \end{align*}\] Recall that for a random variable \(X\), its variance is \(V(X)=E\left(X^2\right)-\left(E(X)\right)^2\). Since \(E(X)=0\), you have \(V(X)=E\left(X^2\right)-0=1\). \[\begin{align*}E\left[(X+1)(X-3Y)\right]&=E\left[X^2\right]+0-3E\left[XY\right]-3(0)\\ &=1-3E\left[XY\right] \end{align*}\] Now let's use what you know about covariance. \[\begin{align*}\text{Covar}(X,Y)&=E\bigg[(X-E(X))~(Y-E(Y))\bigg]\\ &=E\bigg[XY-E(X)Y-E(Y)X+E(X)E(Y)\bigg]\\ &=E(XY)-E\bigg[E(X)Y\bigg]-E\bigg[E(Y)X\bigg]+E\bigg[E(X)E(Y)\bigg]\\ &=E(XY)-E(X)E(Y)-E(X)E(Y)+E(X)E(Y)\\ &=E(XY)-E(X)E(Y) \end{align*}\] Use the given information. The covariance is 1, and the two expected values are both 0: \[1=E(XY)-0\cdot0~~\Rightarrow~~E(XY)=1\] \[E\left[(X+1)(X-3Y)\right]=1-3(1)=-2\]

OpenStudy (anonymous):

awesomework